# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ 6.2 An inspection of the quiet star hosting WASP-39 b
| H2O (base model) | 1.29520.0042 | - | - | 367.330.22 | ||
| H2O + clouds | - | 371.490.23 | ||||
| H2O + clouds + K | 370.890.23 | |||||
| H2O + Stellar activity | - | - | 367.240.25 | |||
| H2O + Stellar activity + clouds | - | 368.490.24 | ||||
| H2O + Stellar activity + clouds + K | 367.590.25 | |||||
@@
| H2O (base model) | - | - | 361.980.24 | |||
| H2O + clouds | - | 372.660.24 | ||||
| H2O + clouds + K | 371.960.24 | |||||
| H2O + Stellar activity | - | - | 364.410.24 | |||
| H2O + Stellar activity + clouds | - | 367.870.27 | ||||
| H2O + Stellar activity + clouds + K | 369.550.26 | |||||
@@
WASP-39 b is a Jupiter-size exoplanet that transits a mature G-type dwarf star (Faedi et al., 2011). Observations using various telescopes have provided insights into WASP-39 b’s atmospheric composition and properties. Using the Hubble Space Telescope and the Very Large Telescope FORS2 instrument, studies led by Fischer et al. (2016) and Nikolov et al. (2016) conducted transmission spectroscopy on WASP-39 b, revealing Rayleigh scattering, sodium and potassium absorption, and a predominantly clear atmosphere, consistent across both ground-based and space-based observations. Further analyses with WFC3 G102 and G141 data exhibited prominent water absorption features in the mid-infrared (Wakeford et al., 2018), while ground-based optical data suggested super-solar atmospheric metallicities (Kirk et al., 2019). Our findings are largely consistent with previous studies. In Case 1 (Fig. 15), we observe a haze-free spectrum with a distinct Rayleigh scattering slope. Contrary to previous analyses (Sing et al., 2016), our best-fit model contains a layer of opaque clouds at Pa, alongside water vapour at a volume mixing ratio of (Table 4). Similarly, Case 2 also prefers the presence of clouds to a clear model, as in Pinhas et al. (2018). However, in this scenario we retrieve a water abundance exceeding solar levels, likely influenced by a potential small offset in the STIS G430L visit 2 data compared to the rest of the observations (Fig. 16). The less pronounced optical slope in this STIS visit prompts the retrieval to increase the water content to counteract the nearly flat spectrum in the optical range. Confirmation of wavelength-independent clouds, recently discovered through JWST NIRSpec/PRISM (Rustamkulov et al., 2023), supports our findings and the absence of sharp alkali signatures in WASP-39 b’s spectrum reinforces this conclusion. Although alkali metals were previously detected by FORS2 and more recently by the James Webb Space Telescope, including in our independent analysis of the WASP-39 b NIRISS transit (Feinstein et al., 2023), the STIS data quality and its limited resolution hinder the identification of these distinct signatures.
@@
There is consensus across previous studies, except from Pinhas et al. (2018) who report substantial evidence for stellar heterogeneities, that the star appears to be magnetically quiet: Sing et al. (2016) report a value of = -4.994, while Mancini et al. (2018) determine an S-index of = 0.1830.017 and a corresponding = -4.970.06. Claudi et al. (2024) report an S-index of = 0.1720.035 and a corresponding = -4.97780.1726, or = -4.7950.2549 when the correction detailed in Mittag et al. (2013) is applied, all of which suggest relative chromospheric inactivity. The Bayesian evidence for most of our active-star scenarios are consistently lower than those that treat the star as being inactive. Specifically, introducing stellar contamination to the cloudy base model in both spectral cases fails to improve the fit. In Figure 4, we observe a Bayes factor of approximately 0 for Case 1 and around 2.5 for Case 2. These values suggest inconclusive evidence for Case 1 and a weak preference for the stellar model in Case 2. This reinforces the fact that stellar activity is likely not affecting the planetary spectrum. The higher evidence for Case 2 may be attributed to a flatter optical region from 0.4 to 0.6 m, which the retrieval attempts to explain with a high concentration of faculae (with a retrieved filling factor of ). However, as mentioned above, the inclusion of clouds further favours the base model over the stellar one, resulting in a Bayes factor of 4.79 in moderate preference of the base model. This serves as another indication supporting the robustness of our stellar contamination model, which appears to correctly identify the absence/non-necessity of activity where it is not playing a critical role.
@@
The SAD for this planet corroborates the Bayes factor results. In Case 2, we identify an SAD of 0.9160.313, the lowest among all Case 2 spectra, while for Case 1, the SAD is the second lowest at 0.8730.403. Both SAD values being less than 1 reaffirm the star’s inactivity. Additionally, WASP-39 b exhibits a low SAT of 2.679%1.152%, placing it among the planets that do not display strong spectral modulations across the two observational epochs available to us. Notably, no clear spot- or faculae-crossing events are evident in the STIS white light curves, further supporting the star’s quiet nature. The absence of lithium in the star’s spectrum, the relatively slow rotation rate of 44.03.9 days (Mancini et al., 2018), and the lack of stellar activity, as evidenced by the absence of Caii H & K emission, collectively suggest that the host star is relatively old, consistent with the age of 7.21.1 Gyrs determined in Mancini et al. (2018).
@@ 6.3 A notoriously active star observed whilst sleeping: the curious apparent quiescence of HAT-P-11
Observations of HAT-P-11 b’s transmission spectrum using data from the HST, Spitzer and Giant Meterwave Radio Telescope (GMRT) telescopes have shown water vapour absorption (Fraine et al., 2014), helium (Mansfield et al., 2018; Allart et al., 2018), and hints of radio emission (Lecavelier des Etangs et al., 2013; Hodosán et al., 2016). The planet’s atmosphere has been found to be predominantly flat in the optical and with weakened features in the NIR, suggesting a high-altitude cloud layer (Chachan et al., 2019). Our findings (Table 5) are consistent with a cloudy atmosphere in both spectral scenarios (Fig. 17 & 18), that attenuates the water feature at 1.4 m, and masks the potassium feature entirely. Our retrievals detect a faint sodium signature, also attenuated by the presence of a high-altitude cloud layer. The cloudiness in HAT-P-11 b’s atmosphere is supported by various studies indicating that Neptune-like planets are predominantly cloudy (Gao et al., 2020; Brande et al., 2024).
@@
| H2O (base model) | - | - | 412.030.22 | |||
| H2O + Na | - | 412.470.22 | ||||
| H2O + clouds + Na | 416.660.24 | |||||
| H2O (STIS shifted down) | - | - | 415.100.22 | |||
| H2O + Stellar activity | - | - | 409.640.24 | |||
| H2O + Stellar activity + Na | - | 409.700.24 | ||||
| H2O + Stellar activity + clouds + Na | 413.820.26 | |||||
@@
| H2O (base model) | 0.42230.0009 | - | - | 383.180.23 | ||
| H2O + Na | - | 384.210.23 | ||||
| H2O + clouds + Na | 389.500.24 | |||||
| H2O (STIS shifted down) | - | - | 399.800.22 | |||
| H2O + Stellar activity | - | - | 388.250.25 | |||
| H2O + Stellar activity + Na | 0.41480.0022 | - | 387.720.25 | |||
| H2O + Stellar activity + clouds + Na | 388.900.26 | |||||
@@
Numerous studies have reported HAT-P-11 to be one of the most active exoplanet hosts. The star exhibits a high level of chromospheric activity, as evidenced by observations of the Caii H & K lines (Morris et al., 2017b). The reported activity indexes are consistent with a high level of activity for HAT-P-11, with a S-index () equal to 0.580 and a = -4.567 respectively (Knutson et al., 2010). More recently, values of = 0.5440.014 and = -4.68080.081 or = -4.30170.009 when the correction detailed in Mittag et al. (2013) is applied are obtained in Claudi et al. (2024). Kepler transit data further reveals the occurrence of frequent star spot crossing events, concentrated primarily at phases -0.002 and +0.006, indicating a transit path nearly perpendicular to the stellar equator (Deming et al., 2011). Studies by Holczer et al. (2015) find that spots on HAT-P-11 are predominantly located at low stellar latitudes. The distribution of spots on HAT-P-11 mirrors that of sunspots near solar activity maximum, with a mean spot latitude of approximately 16 degrees (Morris et al., 2017a). Moreover, the majority of starspots on HAT-P-11 closely resemble the sizes of sunspots at solar maximum, with an estimated mean spotted area coverage two orders of magnitude greater than typical solar spotted areas (Morris et al., 2017a). Observations acquired with the Gran Telescopio Canarias suggest a spot filling factor of = and a spot-to-photosphere temperature difference of T = K (Murgas et al., 2019). However, our retrieval results do not present compelling evidence for stellar activity contamination. In Case 1, the Bayes Factor indicates a weak preference for the base model, supported by a SAD value of 0.9980.286. For Case 2, our retrievals report a stronger inclination towards stellar contamination, with a slightly above 5. Nevertheless, a SAD of 1.2360.384 still suggests a weak preference for stellar activity, the possibility of a quiet star remaining plausible considering the measurement uncertainty of the metric. The system’s high temporal variability is highlighted by its Stellar Activity Temporal (SAT) value, which exceeds 9%. The exoplanet was observed by the STIS instrument at three different times over the course of three months, namely on 22 February 2017, 12 April 2017 and 26 May 2017. The substantial variations among the optical data are unlikely to arise from photospheric activity within this brief 3-month period. This is especially noteworthy as the star exhibits an unexpectedly long activity cycle, known to last for over 10 years. Moreover, its unusually strong chromospheric emission, compared to planet-hosting stars of similar effective temperature and rotation period, may be influenced by tides raised by its planetary companion (Morris et al., 2017b). What we observed is likely due to the interplay between an active photosphere, a disruptive flaring event and perhaps more complex cloud parameterisations such as hazes made of tholins or soots. Our stellar activity characterisation is also likely hindered to an extent by the WFC3 observations, particularly that of G102 which exhibit a downward trend caused by the reduced throughput at the grism’s edge. This effect, coupled with small error bars achievable with the combination of seven different G102 observations, causes the Bayesian approach of the retrieval to prioritise WFC3 obtaining the best possible fit at these longer wavelengths. This may prevent the retrieval from extracting the full information content afforded to us by the STIS observations, where most of the information surrounding stellar contamination lies.
@@ 7 Conclusions
Our findings demonstrate the significant role that stellar contamination may have in all exoplanet spectra observations: therefore, comprehending, modelling, and correcting for the impact of stellar activity is important for a complete characterisation of exoplanet atmospheres. This issue becomes particularly relevant with the ongoing observations of small planets, which are statistically more likely to orbit active stars, performed by JWST. Additionally, addressing stellar activity becomes even more pertinent in anticipation of the extensive chemical surveys to be conducted by space telescopes such as Ariel and Twinkle in the coming decade. For these reasons, we argue that observational campaigns aimed at thoroughly investigating the activity of exoplanets’ stellar hosts through time would significantly contribute to the success of the exoplanet field.
@@ Appendix B Planetary Parameters
| WASP-31 b | 0.4780.029 | 1.5490.050 | 3.40590960.000005 | 84.410.22 | 8.0020.025 ∗ | 0.0 | - | Anderson et al. (2011) | |
| WASP-39 b | 0.280.03 | 1.270.04 | 4.0552590.000009 | 11.6770.027∗ | 0.0 | - | Faedi et al. (2011) | ||
@@
| WASP-69 b | 0.2600.017 | 1.0570.047 | 3.86813820.0000017 | 86.710.20 | 11.9680.036∗ | 0.0 | - | Anderson et al. (2014) | |
| WASP-74 b | 0.9500.060 | 1.560.06 | 2.1377500.000001 | 79.810.24 | 4.8510.041∗ | 0.0 | - | Hellier et al. (2015) | |

### v3

@@ 6.2 An inspection of the quiet star hosting WASP-39 b
H2O (base model) | - | - | 367.30.2 | |||
H2O + clouds | - | 371.50.2 | ||||
H2O + clouds + K | 370.90.2 | |||||
H2O + Stellar activity | - | - | 367.20.3 | |||
H2O + Stellar activity + clouds | - | 368.50.2 | ||||
H2O + Stellar activity + clouds + K | 367.60.3 | |||||
@@
H2O (base model) | - | - | 362.00.2 | |||
H2O + clouds | - | 372.70.2 | ||||
H2O + clouds + K | 372.00.2 | |||||
H2O + Stellar activity | - | - | 364.40.2 | |||
H2O + Stellar activity + clouds | - | 367.90.3 | ||||
H2O + Stellar activity + clouds + K | 369.60.3 |
@@
WASP-39 b is a Jupiter-size exoplanet that transits a mature G-type dwarf star (Faedi et al., 2011). Observations using various telescopes have provided insights into WASP-39 b’s atmospheric composition and properties. Using the Hubble Space Telescope and the Very Large Telescope FORS2 instrument, studies led by Fischer et al. (2016) and Nikolov et al. (2016) conducted transmission spectroscopy on WASP-39 b. They revealed Rayleigh scattering, sodium and potassium absorption, and a predominantly clear atmosphere, consistent across both ground-based and space-based observations. Further analyses with WFC3 G102 and G141 data exhibited prominent water absorption features in the mid-infrared (Wakeford et al., 2018), while ground-based optical data suggested super-solar atmospheric metallicities (Kirk et al., 2019). Our findings are largely consistent with previous studies. In Case 1 (Fig. 16), we observe a haze-free spectrum with a distinct Rayleigh scattering slope. Contrary to previous analyses (Sing et al., 2016), our best-fit model contains a layer of opaque clouds at Pa, alongside water vapour at a volume mixing ratio of (Table 3). Similarly, Case 2 also prefers the presence of clouds to a clear model, as in Pinhas et al. (2018). However, in this scenario we retrieve a water abundance exceeding solar levels, likely influenced by a potential small offset in the STIS G430L visit 2 data compared to the rest of the observations (Fig. 17). The less pronounced optical slope in this STIS visit prompts the retrieval to increase the water content to counteract the nearly flat spectrum in the optical range. Confirmation of wavelength-independent clouds, recently discovered through JWST NIRSpec/PRISM (Rustamkulov et al., 2023), supports our findings, and the absence of sharp alkali signatures in WASP-39 b’s spectrum reinforces this conclusion. Although alkali metals were previously detected by FORS2 and more recently by the James Webb Space Telescope, including in our independent analysis of the WASP-39 b NIRISS transit (Feinstein et al., 2023), the STIS data quality and its limited resolution hinder the identification of these distinct signatures.
@@
There is consensus across previous studies, except from Pinhas et al. (2018) who report substantial evidence for stellar heterogeneities, that the star appears to be magnetically quiet: Sing et al. (2016) report a value of = -4.994, while Mancini et al. (2018) determine an S-index of = 0.180.02 and a corresponding = -4.970.06. Claudi et al. (2024) report an S-index of = 0.170.04 and a corresponding = -5.00.2, or = -4.80.6 when the correction detailed in Mittag et al. (2013) is applied, all of which suggest relative chromospheric inactivity. The Bayesian evidence for most of our active-star scenarios are consistently lower than those that treat the star as being inactive. Specifically, introducing stellar contamination to the cloudy base model in both spectral cases fails to improve the fit. In Figure 4, we observe a Bayes factor of approximately 0 for Case 1 and around 2.5 for Case 2. These values suggest inconclusive evidence for Case 1 and a weak preference for the stellar model in Case 2. This reinforces the fact that stellar activity is likely not affecting the planetary spectrum. The higher evidence for Case 2 may be attributed to a flatter optical region from 0.4 to 0.6 m, which the retrieval attempts to explain with a high concentration of faculae (with a retrieved filling factor of ). However, as mentioned above, the inclusion of clouds further favours the base model over the stellar one, resulting in a Bayes factor of 4.79 in moderate preference of the base model. This serves as another indication supporting the robustness of our stellar contamination model, which appears to correctly identify the absence/non-necessity of activity where it is not playing a critical role.
@@
The SAD for this planet corroborates the Bayes factor results. In Case 2, we identify an SAD of 0.90.3, the lowest among all Case 2 spectra, while for Case 1, the SAD is the second lowest at 0.90.4. Both SAD values being less than 1 reaffirm the star’s inactivity. Additionally, WASP-39 b exhibits a low SAT of 31%, placing it among the planets that do not display strong spectral modulations across the two observational epochs available to us. Notably, no clear spot- or faculae-crossing events are evident in the STIS white light curves, further supporting the star’s quiet nature. The absence of lithium in the star’s spectrum, the relatively slow rotation rate of 444 days (Mancini et al., 2018), and the lack of stellar activity, as evidenced by the absence of Caii H & K emission, collectively suggest that the host star is relatively old, consistent with the age of 71 Gyrs determined in Mancini et al. (2018).
@@ 6.3 A notoriously active star observed whilst sleeping: the curious apparent quiescence of HAT-P-11
Observations of HAT-P-11 b’s transmission spectrum using data from the HST, Spitzer and Giant Meterwave Radio Telescope (GMRT) telescopes have shown water vapour absorption (Fraine et al., 2014), helium (Mansfield et al., 2018; Allart et al., 2018), and hints of radio emission (Lecavelier des Etangs et al., 2013; Hodosán et al., 2016). The planet’s atmosphere has been found to be predominantly flat in the optical and with weakened features in the NIR, suggesting a high-altitude cloud layer (Chachan et al., 2019). Our findings (Table 4) are consistent with a cloudy atmosphere in both spectral scenarios (Fig. 18 & 19), that attenuates the water feature at 1.4 m, and masks the potassium feature entirely. Our retrievals detect a faint sodium signature, also attenuated by the presence of a high-altitude cloud layer. The cloudiness in HAT-P-11 b’s atmosphere is supported by various studies indicating that Neptune-like planets are predominantly cloudy (Gao et al., 2020; Brande et al., 2024).
@@
H2O (base model) | - | - | 412.00.2 | |||
H2O + Na | - | 412.50.2 | ||||
H2O + clouds + Na | 416.70.2 | |||||
H2O (STIS shifted down) | - | - | 415.10.2 | |||
H2O + Stellar activity | - | - | 409.60.2 | |||
H2O + Stellar activity + Na | - | 409.70.2 | ||||
H2O + Stellar activity + clouds + Na | 413.80.3 | |||||
@@
H2O (base model) | - | - | 383.20.2 | |||
H2O + Na | - | 384.20.2 | ||||
H2O + clouds + Na | 389.50.2 | |||||
H2O (STIS shifted down) | - | - | 399.80.2 | |||
H2O + Stellar activity | - | - | 388.30.3 | |||
H2O + Stellar activity + Na | - | 387.70.3 | ||||
H2O + Stellar activity + clouds + Na | 388.90.3 |
@@
Numerous studies have reported HAT-P-11 to be one of the most active exoplanet hosts. The star exhibits a high level of chromospheric activity, as evidenced by observations of the Caii H & K lines (Morris et al., 2017b). The reported activity indexes are consistent with a high level of activity for HAT-P-11, with a S-index () equal to 0.580 and a = -4.567 respectively (Knutson et al., 2010). More recently, values of = 0.540.01 and = -4.680.08 or = -4.3020.009 when the correction detailed in Mittag et al. (2013) is applied are obtained in Claudi et al. (2024). Kepler transit data further reveals the occurrence of frequent star spot crossing events, concentrated primarily at phases -0.002 and +0.006, indicating a transit path nearly perpendicular to the stellar equator (Deming et al., 2011). Studies by Holczer et al. (2015) find that spots on HAT-P-11 are predominantly located at low stellar latitudes. The distribution of spots on HAT-P-11 mirrors that of sunspots near solar activity maximum, with a mean spot latitude of approximately 16 degrees (Morris et al., 2017a). Moreover, the majority of starspots on HAT-P-11 closely resemble the sizes of sunspots at solar maximum, with an estimated mean spotted area coverage two orders of magnitude greater than typical solar spotted areas (Morris et al., 2017a). Observations acquired with the Gran Telescopio Canarias suggest a spot filling factor of = and a spot-to-photosphere temperature difference of T = K (Murgas et al., 2019). However, our retrieval results do not present compelling evidence for stellar activity contamination. In Case 1, the Bayes factor indicates a weak preference for the base model, supported by a SAD value of 1.00.3. For Case 2, our retrievals report a stronger inclination towards stellar contamination, with a slightly above 5. Nevertheless, a SAD of 1.20.4 still suggests a weak preference for stellar activity, the possibility of a quiet star remaining plausible considering the measurement uncertainty of the metric. The system’s high temporal variability is highlighted by its Stellar Activity Temporal (SAT) value, which exceeds 9%. The exoplanet was observed by the STIS instrument at three different times over the course of three months, namely on 22 February 2017, 12 April 2017 and 26 May 2017. The substantial variations among the optical data are unlikely to arise from photospheric activity within this brief 3-month period. This is especially noteworthy as the star exhibits an unexpectedly long activity cycle, known to last for over 10 years. Moreover, its unusually strong chromospheric emission, compared to planet-hosting stars of similar effective temperature and rotation period, may be influenced by tides raised by its planetary companion (Morris et al., 2017b). What we observed is likely due to the interplay between an active photosphere, a disruptive flaring event and perhaps more complex cloud parameterisations such as hazes made of tholins or soots. Our stellar activity characterisation is also likely hindered to an extent by the WFC3 observations, particularly that of G102 which exhibit a downward trend caused by the reduced throughput at the grism’s edge. This effect, coupled with small error bars achievable with the combination of seven different G102 observations, causes the Bayesian approach of the retrieval to prioritise WFC3 obtaining the best possible fit at these longer wavelengths. This may prevent the retrieval from extracting the full information content afforded to us by the STIS observations, where most of the information surrounding stellar contamination lies.
@@ 7 Conclusions
Our findings demonstrate the significant role that stellar contamination may have in all exoplanet spectra observations. Therefore, comprehending, modelling, and correcting for the impact of stellar activity is important for a complete characterisation of exoplanet atmospheres. This issue becomes particularly relevant with the ongoing observations of small planets, which are statistically more likely to orbit active stars, performed by JWST. Additionally, addressing stellar activity becomes even more pertinent in anticipation of the extensive chemical surveys to be conducted by space telescopes such as Ariel and Twinkle in the coming decade.
Ideally, the community should aim to have simultaneous photometry alongside all of the future exoplanetary observations. This is to place exoplanets in the correct stellar context of both localised photospheric activity and proximity to stellar minima or maxima. As stellar variability operates over a wide range of timescales, two different observing strategies would be recommended. The first would involve a finely sampled, short cadence survey to capture the stellar rotational modulation, typically on the order of days to weeks. The second should use a longer cadence survey conducted over many years to constrain long-term magnetic cycles analogous to the 11-year solar cycle.
For both approaches we would recommend the use of continuous photospheric monitoring, preferably in the B or V bands where stellar activity effects are the most enhanced. Existing ground-based networks of moderately-sized telescopes ( 40 cm) such as the Las Cumbres Observatory (LCO) or those used by the MEarth survey could be ideal for this purpose. In the meantime, robust homogeneous analyses of TESS light curves will provide invaluable insights into a star’s activity levels. Particularly for the targets that have already been observed over multiple epochs, we could extrapolate activity information on the timescale of their rotation period and potentially on an epoch-to-epoch basis.
Spectroscopic observations of e.g. Ca H&K lines would also be advantageous. However, the chromospheric activity measured may not be a direct representation of the photospheric contamination observed. Equally, the time and resource intensive nature of such observations may make this approach less plausible for the large homogeneous survey we recommend.
Additionally, we emphasise the importance of data reduction homogeneity, to prevent the introduction of spurious signals through systematics and/or calibration errors. This could perhaps make the exploitation of pre-existing archival data difficult but not impossible.
The observational strategies to monitor stellar activity highlighted above would significantly contribute to the success of the exoplanet field.
@@ Acknowledgements
We would like to express our gratitude to the anonymous reviewer for their invaluable feedback, which significantly enhanced the quality of our work.
@@
The observations analysed in this study can be accessed viahttps://doi.org/10.17909/7nt2-mk29 (catalog 10.17909/7nt2-mk29).
@@ Appendix B Planetary Parameters
WASP-31 b | 0.4780.029 | 1.5490.050 | 3.40590960.000005 | 84.410.22 | 8.000.03 ∗ | 0.0 | - | Anderson et al. (2011) | |
WASP-39 b | 0.280.03 | 1.270.04 | 4.0552590.000009 | 11.680.03 ∗ | 0.0 | - | Faedi et al. (2011) | ||
@@
WASP-69 b | 0.2600.017 | 1.0570.047 | 3.86813820.0000017 | 86.710.20 | 11.970.04 ∗ | 0.0 | - | Anderson et al. (2014) | |
WASP-74 b | 0.9500.060 | 1.560.06 | 2.1377500.000001 | 79.810.24 | 4.850.04 ∗ | 0.0 | - | Hellier et al. (2015) | |